In a simple​ economy, the savings function is Upper S equals negative 200 plus 0.20 Upper Y. Investment is equal to 500. In this​ economy, equilibrium GDP is ​$ nothing. ​(Round your answer the nearest​ dollar.)  
Answer #1

In a simple​ economy,

S=-200+0.20Y
I=500

In equilibrium

S=I

-200+0.20Y=500

0.20Y=500+200

Y=700/0.20

=3,500

the savings function is Upper S equals negative 200 plus 0.20 Upper Y. Investment is equal to 500. In this​ economy, equilibrium GDP is ​$ nothing. ​(Round your answer the nearest​ dollar.)